OpenStudy (anonymous):

Find the values of k that make 2x^2 + kx + 5 factorable.

OpenStudy (ksaimouli):

use quadratic formula

OpenStudy (anonymous):

@ksaimouli wtf is that

OpenStudy (ksaimouli):

okay not introduced dont worry

OpenStudy (ksaimouli):

what is 2*5

OpenStudy (anonymous):

10

OpenStudy (ksaimouli):

yup so what r some multiples

OpenStudy (ksaimouli):

of 10

OpenStudy (anonymous):

Well I know like how to factor an polynomial like that but not how to find a variable in one

OpenStudy (ksaimouli):

lets say these for now 5 and 2

OpenStudy (ksaimouli):

add 5+2=7 so the variable is 7

OpenStudy (anonymous):

wait so then k would be 7

OpenStudy (ksaimouli):

yup

jimthompson5910 (jim_thompson5910):

there are 3 more possible answers, so keep that in mind

OpenStudy (ksaimouli):

that if one of them

OpenStudy (ksaimouli):

if->is

OpenStudy (anonymous):

So it would be 7 or 11

jimthompson5910 (jim_thompson5910):

good, then -7 and -11 are the other two because 2x^2 - 7x + 5 and 2x^2 - 11x + 5 are both factorable

jimthompson5910 (jim_thompson5910):

since -2 + (-5) = -7 and -1 + (-10) = -11